# Graph neural networks for learning liquid simulations in dynamic scenes containing kinematic objects

### v1

@@ 4.1 Ground Truth Simulator
We use NVIDIA Isaac Sim [38] as our ground truth simulator. Isaac Sim is a robotics simulation platform that also supports particle simulation based on the Position Based Dynamics (PBD) [39] method. To investigate the ability of our system to generalize from a small set of training data, our primary dataset consisted of a scene with a collection of liquid particles and three rigid bodies: a jug (kinematic), a cup (stationary), and a floor (stationary). We adjusted the properties of particles to approximate those of water. Each simulation episode was initialized as follows: the cup was placed at a random location, the floor was positioned underneath it, and the jug was placed above the floor, containing the liquid particles. To enable the network to learn liquid dynamics under a wide range of jug motions, as well as the effects of liquid collisions with stationary objects (cup and floor), we create three types of simulation scenarios:
@@ 4.2 Evaluation
Baseline models We compare our model against recent GNN-based particle simulators that support kinematic rigid body manipulation, specifically Graph Network Simulator (GNS) [10] and MeshGraphNet [11]. Both models handle liquid-object interaction using node-based collision. GNS uses a single edge set to model both liquid-liquid and liquid-object interactions. In contrast, MeshGraphNet separates liquid-liquid and liquid-object interactions into two distinct edge sets. We evaluate two variants of MeshGraphNet based on how the liquid-object edge set is defined. In the first variant (MGN), it includes only liquid-object edges. In the second variant (MGN*), mesh connectivity edges are included alongside the liquid-object edges. All 3 models use a single node set to represent both particle and mesh nodes. As a result, the 6-DOF control inputs for manipulating a rigid body are incorporated into the node features of all the nodes in the graph. This implicit representation of control inputs limits these models to manipulating only one rigid object per simulation. We use the same neighborhood radii as in our model: $r_{l}$ for liquid-liquid collision and $r_{ol}$ for liquid-object collision. See Appendix A.2 for complete details of the baseline models.
@@ 4.3 Results
Generalization A key advantage of our model is its ability to generalize to novel object shapes. Figure 4 compares rollouts from our Full-model, GNS, and MGN on the Wine-Beaker and Martini-Bottle test sets. Despite never encountering these shapes during training, our model accurately simulates the liquid dynamics, closely matching the ground truth behavior. Importantly, our Ablated-model also performs comparably well, suggesting that both models effectively learn to represent local surface geometry through our liquid-object collision handling. In contrast, the GNS and MGN baselines exhibit substantial limitations. With the Beaker, GNS shows pronounced tunneling effects, with particles incorrectly passing through the spout walls. While MGN is able to channel particles through the spout, it still suffers from tunneling toward the end of the rollout. With the WineBottle, both GNS and MGN struggle to simulate the liquid flow as the particles approach the narrow neck, resulting in significant particle tunneling. This difficulty stems from their reliance on mesh vertices alone to handle liquid-object collisions, which fails to capture finer geometric details of these novel shapes. Importantly, this is despite increasing the mesh vertex density of the objects by 2-5 times to facilitate better shape information. MGN* performs worse than both GNS and MGN across all evaluated scenarios. Table 1 summarizes the quantitative results for each model, while Figure 6 shows the step-wise error over the trajectory. Our Full-model consistently achieves the best scores across most test scenarios. Additional snapshots can be found in Appendix A.3, and videos of simulations can be found in the supplementary material.
@@
We further validated our Full-model’s generalization capabilities by simulating liquid pouring from the classic Utah teapot, which has intricate geometry. As shown in Figure 1, the Full-model (chamfer loss of 0.078) successfully simulates realistic liquid dynamics even for this challenging shape. The Ablated-model (chamfer loss of 0.301) however, exhibited some leakage with the Utah teapot, suggesting a potential limitation in handling highly complex geometries. The difference in performance may be due to the unavailability of object node $\mathcal{V}^{O}$ updates during the message-passing process in the Ablated-model. Further investigation, including exploring longer training schedules for the Ablated-model, would be needed to fully understand this behavior.
@@ 4.5 Control on learned dynamics
A key advantage of a fully differentiable environment is the ability to solve tasks using gradient-based optimization techniques. To showcase the utility of our model, we designed a pouring task where the objective is to pour a specified percentage of the cup’s total volume from the jug into the cup. In the initial state, the jug is placed in an upright position next to the cup, such that rotating the jug around the $x$-axis will cause liquid to pour into the cup. We use an MPC controller with the control variable being the jug’s rotational acceleration angle around the $x$-axis. The control inputs are initialized randomly. Following a task formulation similar to [27], we implement a two-stage cost function optimized using Adam [40]. In the first stage, the cost is the $L2$ distance between all liquid particles and a target point on the rim of the jug closest to the cup. This encourages the jug to rotate towards the cup to let the liquid flow into the cup. Once the target fill level is reached, the second stage activates. This stage uses a regularization term on the jug’s rotation angle, encouraging it to return to its initial upright position.
@@ 5 Discussion
While promising, our current framework is limited in its ability to simulate fluids with diverse physical properties, such as viscosity. Future work will focus on incorporating richer representations of material properties within the network architecture. Additionally, our framework is currently designed exclusively for learning fluid dynamics and, as such, cannot handle rigid body dynamics when interacting with particles. For instance, it cannot simulate a scenario where flowing water displaces or carries a rigid object. Extending the architecture to incorporate learning rigid body simulation is a natural next step toward realising a general-purpose learned simulator that can handle every complex physical scenario.
@@ A.1 Architecture and training
All the MLPs in our architecture contain 2 hidden layers of 128 units each with LayerNorm [41] and an output layer of 128 units. Only in the case of the decoder, the output layer has 3 units to decode the 3D acceleration of particles. We chose $P=10$ as the number of processor blocks.
It has been shown that perturbing the inputs with noise makes the GNN models more robust to noisy inputs [10]. This is particularly helpful when generating rollouts where the network is fed with its own noisy, previous predictions as input. We choose a noise value $\mathcal{N}(0,\sigma=0.00067)$. Additionally, all the inputs and outputs were normalised.
We use Adam [40] to optimize the $L2$ loss on one-step acceleration predictions of the liquid particles with an exponential decaying learning rate from 1e-4 to 1e-6. The predicted particle accelerations are then integrated to produce the next-step positions using the finite-differences method.
The models were trained on a single Nvidia 4090 GPU up to a maximum of $2M$ gradient steps with a batch size of 20. The training time took about 10 days.
@@ A.2 Baseline model details
GNS baseline has a graph of the form $\mathcal{G}=(\mathcal{V},\mathcal{E})$, while MeshGraphNet (MGN and MGN*) models use a multigraph of the form $\mathcal{G}=(\mathcal{V},\mathcal{E}^{L},\mathcal{E}^{M})$. For all three baselines, the node set $\mathcal{V}$ contains both liquid particles and rigid bodies. To represent a rigid body, its mesh vertices are used as graph nodes. In order to distinguish between particle nodes and mesh nodes, one-hot encoding is incorporated into the node features. Unlike our model, GNS does not explicitly represent the floor as a separate object; instead, the distance from each particle to the floor is included as part of each particle’s node features. We use the same floor representation for MeshGraphNets as well.
@@
Since collision with the particles is handled at the mesh vertex level in the baseline models, a low vertex count negatively impacts performance. To ensure a fair comparison, we oversampled all objects used in the baseline evaluations. Table 2 summarizes the number of nodes and edges used in both our models and the baseline models.
@@
A.3 Additional Figures

### v2

@@ 4.1 Ground Truth Simulator
We use NVIDIA Isaac Sim [isaac_sim] as our ground truth simulator. Isaac Sim is a robotics simulation platform that also supports particle simulation based on the Position Based Dynamics (PBD) [pbd] method. To investigate the ability of our system to generalize from a small set of training data, our primary dataset consisted of a scene with a collection of liquid particles and three rigid bodies: a jug (kinematic), a cup (stationary), and a floor (stationary). We adjusted the properties of particles to approximate those of water. Each simulation episode was initialized as follows: the cup was placed at a random location, the floor was positioned underneath it, and the jug was placed above the floor, containing the liquid particles. To enable the network to learn liquid dynamics under a wide range of jug motions, as well as the effects of liquid collisions with stationary objects (cup and floor), we create three types of simulation scenarios:
@@ 4.2 Evaluation
Baseline models We compare our model against recent GNN-based particle simulators that support kinematic rigid body manipulation, specifically Graph Network Simulator (GNS) [sanchez2020learning] and MeshGraphNet [meshgraphnet]. Both models handle liquid-object interaction using node-based collision. GNS uses a single edge set to model both liquid-liquid and liquid-object interactions. In contrast, MeshGraphNet separates liquid-liquid and liquid-object interactions into two distinct edge sets. We evaluate two variants of MeshGraphNet based on how the liquid-object edge set is defined. In the first variant (MGN), it includes only liquid-object edges. In the second variant (MGN*), mesh connectivity edges are included alongside the liquid-object edges. All 3 models use a single node set to represent both particle and mesh nodes. As a result, the 6-DOF control inputs for manipulating a rigid body are incorporated into the node features of all the nodes in the graph. This implicit representation of control inputs limits these models to manipulating only one rigid object per simulation. We use the same neighborhood radii as in our model: $r_{l}$ for liquid-liquid collision and $r_{ol}$ for liquid-object collision. See Appendix A.2 for complete details of the baseline models.
@@ 4.3 Results
Generalization A key advantage of our model is its ability to generalize to novel object shapes. Figure 4 compares rollouts from our Full-model, GNS, and MGN on the Wine-Beaker and Martini-Bottle test sets. Despite never encountering these shapes during training, our model accurately simulates the liquid dynamics, closely matching the ground truth behavior. Importantly, our Ablated-model also performs comparably well, suggesting that both models effectively learn to represent local surface geometry through our liquid-object collision handling. In contrast, the GNS and MGN baselines exhibit substantial limitations. With the Beaker, GNS shows pronounced tunneling effects, with particles incorrectly passing through the spout walls. While MGN is able to channel particles through the spout, it still suffers from tunneling toward the end of the rollout. With the WineBottle, both GNS and MGN struggle to simulate the liquid flow as the particles approach the narrow neck, resulting in significant particle tunneling. This difficulty stems from their reliance on mesh vertices alone to handle liquid-object collisions, which fails to capture finer geometric details of these novel shapes. Importantly, this is despite increasing the mesh vertex density of the objects by 2-5 times to facilitate better shape information. MGN* performs worse than both GNS and MGN across all evaluated scenarios. Table 1 summarizes the quantitative results for each model, while Figure 6 shows the step-wise error over the trajectory. Our Full-model consistently achieves the best scores across most test scenarios. Additional snapshots can be found in Appendix A.4.
@@
We further validated the generalization capabilities of our Full-model by simulating liquid pouring from the classic Utah teapot, which features intricate geometry. As shown in Figure 1, the Full-model (chamfer loss of 0.078) successfully simulates realistic liquid dynamics even for this challenging shape. The Ablated-model (chamfer loss of 0.301), however, exhibited some leakage with the Utah teapot, suggesting a potential limitation in handling highly complex geometries. The difference in performance may be due to the unavailability of object node $\mathcal{V}^{O}$ updates during the message-passing process in the Ablated-model.
In practice, since both the Full-model and Ablated-model offer similar performance in most scenarios, the two can be seen as complementary. The Ablated-model is suitable for environments with rigid objects of relatively simple shapes, as it reduces computational overhead and memory footprint, while the Full-model is preferable for scenes containing highly intricate geometries.
@@ 4.5 Control on learned dynamics
A key advantage of a fully differentiable environment is the ability to solve tasks using gradient-based optimization techniques. To showcase the utility of our model, we designed a pouring task where the objective is to pour a specified percentage of the cup’s total volume from the jug into the cup. In the initial state, the jug is placed in an upright position next to the cup, such that rotating the jug around the $x$-axis will cause liquid to pour into the cup. We use an MPC controller with the control variable being the jug’s rotational acceleration angle around the $x$-axis. The control inputs are initialized randomly. Following a task formulation similar to [spnet], we implement a two-stage cost function optimized using Adam [ADAM]. In the first stage, the cost is the $L2$ distance between all liquid particles and a target point on the rim of the jug closest to the cup. This encourages the jug to rotate towards the cup to let the liquid flow into the cup. Once the target fill level is reached, the second stage activates. This stage uses a regularization term on the jug’s rotation angle, encouraging it to return to its initial upright position.
@@
4.6 Computational Complexity and Inference Times
For liquid-object collisions, our BVH-based handling reduces complexity compared to node-based baselines. BVH is relatively inexpensive, with a complexity of $\mathcal{O}(\log(K_{i}))$ per query, where $K_{i}$ is the number of vertices in the mesh. In our method, each particle is connected to at most one edge per object, yielding an $\mathcal{O}(NQ)$ complexity for $N$ liquid particles and $Q$ objects. In node-based collision, such as in GNS, a particle may connect to many vertices, resulting in $\mathcal{O}(NK)$ complexity, where $K$ is the total number of object mesh vertices. Since the number of objects is significantly lower than the total number of node vertices($Q<<K$), a substantial speedup is possible.
Table 2: Comparison of Inference Times across all models.
Method Time per step (ms)
Full-model 25.05
Ablated-model 21.34
GNS 75.51
MGN 50.32
MGN* 49.58
Isaac-Sim (equiv.) 10.20
Figure 8: Inference time and memory consumption plotted against the number of particles for our model.
We calculated the rollout inference times per step ($dt=1/60$) for all our models and baselines on the test set (evaluated on a workstation with Nvidia 4090 GPU and 32-core CPU) and report them in Table 2. Indeed, the performance is twice as fast as the fastest baseline methods. However, it is still slower than Isaac-sim. This gap arises from CPU-bound neighborhood computations, a known limitation in these GNN-based methods, as noted in the GNS (supplementary).
For the overall system, our method scales linearly with the number of particles $\mathcal{O}(N)$. We tested this on the stirring environment by varying the number of particles between 2k and 16k particles and report the memory consumption and inference time per step in Figure 8. The results confirm that our model scales linearly with the number of particles.
@@ 5 Discussion
While promising, our current framework is limited in its ability to simulate fluids with diverse physical properties, such as viscosity. Future work will focus on incorporating richer representations of material properties within the network architecture. Additionally, our framework is currently designed exclusively for learning fluid dynamics and, as such, cannot handle rigid body dynamics when interacting with particles. For instance, it cannot simulate a scenario where flowing water displaces or carries a rigid object. However, given our modular graph definition, it is possible to extend our framework to handle this two-way coupling. By introducing trainable non-kinematic rigid-body nodes ($\mathcal{V}^{R}$) alongside liquid nodes ($\mathcal{V}^{L}$), and defining explicit liquid–object ($\mathcal{E}^{LO}$) and object–object ($\mathcal{E}^{OO}$) edges for bidirectional interactions, the framework could learn coupled fluid–rigid dynamics. We consider this a natural extension for future work.
Acknowledgments
This research was supported by the European Research Council (ERC; Consolidator Award ‘ACTOR’-project number ERC-CoG-101045783), by the Hessian Ministry of Higher Education, Research, Science and the Arts with its LOEWE research priority program ‘WhiteBox’, and by ‘The Adaptive Mind’, funded by the Excellence Program of the Hessian Ministry of Higher Education, Science, Research and Art.
@@ A.1 Architecture and training
All the MLPs in our architecture contain 2 hidden layers of 128 units each with LayerNorm [layernorm] and an output layer of 128 units. Only in the case of the decoder, the output layer has 3 units to decode the 3D acceleration of particles. We chose $P=10$ as the number of processor blocks.
It has been shown that perturbing the inputs with noise makes the GNN models more robust to noisy inputs [sanchez2020learning]. This is particularly helpful when generating rollouts where the network is fed with its own noisy, previous predictions as input. We choose a noise value $\mathcal{N}(0,\sigma=0.00067)$. Additionally, all the inputs and outputs were normalised.
We use Adam [ADAM] to optimize the $L2$ loss on one-step acceleration predictions of the liquid particles with an exponentially decaying learning rate from 1e-4 to 1e-6. The predicted particle accelerations are then integrated to produce the next-step positions using the finite-difference method.
The models were trained on a single Nvidia 4090 GPU up to a maximum of $2M$ gradient steps with a batch size of 20. The training period lasted approximately 10 days. The source code and visualizations are available here. [^1] https://github.com/RothkopfLab/fluid-manip-gnn
@@ A.2 Baseline model details
GNS baseline has a graph of the form $\mathcal{G}=(\mathcal{V},\mathcal{E})$, while MeshGraphNet (MGN and MGN*) models use a multigraph of the form $\mathcal{G}=(\mathcal{V},\mathcal{E}^{L},\mathcal{E}^{M})$. For all three baselines, the node set $\mathcal{V}$ contains both liquid particles and rigid bodies. To represent a rigid body, its mesh vertices are used as nodes in a graph. To distinguish between particle nodes and mesh nodes, one-hot encoding is incorporated into the node features. Unlike our model, GNS does not explicitly represent the floor as a separate object; instead, the distance from each particle to the floor is included as part of each particle’s node features. We use the same floor representation for MeshGraphNets as well.
@@
Since collision with particles is handled at the mesh vertex level in the baseline models, a low vertex count negatively impacts performance. To ensure a fair comparison, we oversampled all objects used in the baseline evaluations. Table 3 summarizes the number of nodes and edges used in both our models and the baseline models.
@@
A.3 Failure cases
We tested our model on two scenarios where it fails or performs suboptimally.
Tunneling occurs when object velocities are significantly higher than those seen during training. GNN-based models operate on a fixed time step (dt=1), and hence, if the object pose changes significantly between steps, the object can jump across particles between steps, resulting in missed collisions (“tunneling”). We tested this on the Martini-Vase environment (identical to our training environment) by making the Vase jug rotate rapidly. We observe the tunneling effect at the beginning of the simulation (see video in our project repository). In Classical simulators, this tunneling effect is typically mitigated by using sub-stepping and continuous collision detection (CCD) methods. In learned GNN-based simulators, it is not clear how such remedies can be integrated.
Another case where our model fails is when the liquid motion is caused by shear stress and centrifugal effects. For example, when a cylindrical container with liquid inside spins about its axis, the liquid also rotates along with the container. This behavior cannot be captured by our model, as our liquid–object connectivity is distance-based. When the container rotates around its axis, particle–surface distances remain constant, so the model does not perceive tangential motion. To test this, we created a cement-mixer environment in which liquid particles are placed inside a rotating drum. Our model fails to reproduce the expected rotation of the liquid (see video in our project repository).
A.4 Additional Figures